OpenStudy (jadebirdfly):

Please help i will fan and medal!!! If 3x + 2(1 + x) = 17, what is the value of 6x + 5?

OpenStudy (atrineas):

what does 6x+5=?

OpenStudy (jadebirdfly):

i have no clue

OpenStudy (atrineas):

ok one moment I'll try

OpenStudy (atrineas):

well solving for x on the 3x+2(1+x)=17 I got. . . x=3 Which I'm betting you use to turn the problem 6x+5 into. . . 6(3)+5=?

OpenStudy (jadebirdfly):

23

OpenStudy (atrineas):

correct ^^

OpenStudy (atrineas):

Thats your answer

OpenStudy (jadebirdfly):

thank you!

OpenStudy (atrineas):

No problem ^^